a rectangle has one side equal to(2x+4)cm and a perimeter of (6x+4)cm.what is the area of the rectangle in terms of x?
Answer by mathbird(71)

one side equal to(2x+4)cm
perimeter of (6x+4)cm
formula of perimeter of rectangle=2(length + breadth)
(6x+4)=2{(2x+4)+breadth}
2(3x+2)=2{(2x+4)+breadth}
dividing by 2 on both sides
3x+2=2x+4+breadth
3x+2-2x-4=breadth
x-2=breadth
now;formula of area of rectangle = length *breadth
=(2x+4)*(x-2)
if need more simplification
=>2x^2-4x+4x-8
=2x^2-8